How many milliliters of 0.200M NaOHsolution are needed to react with 28.0 mL of a 0.250 M NiCl2 solution?
Express your answer with the appropriate units.
1
Expert's answer
2015-11-03T12:49:58-0500
NiCl2 (aq) + 2NaOH (aq) = Ni(OH)2 (s) + 2NaCl (aq)
ν (NiCl2)= 0.250 mol/L × 0.028 L = 0.007 mol
ν (NaOH) = 0.007 mol × 2 = 0.014 mol
V (NaOH) = 0.014 mol / 0.200 mol/L = 0.07 L = 70 mL

Answer: 70 milliliters of 0.200M NaOH solution are needed to react with 28.0 mL of a 0.250 M NiCl2 solution.